# Backpropagation-Free Multi-modal On-Device Model Adaptation via Cloud-Device Collaboration

### v2

@@ 2. Related Works
Cloud-Device Collaboration. In the rapidly evolving sphere of cloud-device collaboration, deep learning has become a crucial component, strategically combining the features of cloud-based and on-device machine learning. Although Federated Learning (FL), notably represented by FedAVG (Collins et al., 2022), has traditionally been a significant player, its perceived simplicity has limited its applicability in various practical situations. This dynamic field has seen the rise of innovative methods and techniques designed to optimize the interaction between cloud resources and mobile devices. For instance, MPDA (Yan et al., 2022) skillfully utilizes cloud-based samples to enhance the performance of on-device models. Another interesting approach involves deploying multiple models with similar functionalities, coordinated by a Meta Controller for optimal task allocation, thus expanding the possibilities of collaborative model adaptation. Moreover, DUET (Lv et al., 2023b), inspired by the HyperNetwork concept (Yi et al., 2023; Xian et al., 2021; Zhang et al., 2020; Oscar Chang, 2020), simplifies the device model adaptation process, eliminating the need for on-device training. IDEAL (Lv et al., 2023a) expands parameter generation-based models to the domain of recommender systems, particularly DC-CDR (Zhang et al., 2023). This enhances the generalization capabilities of device recommendation models, though it requires consideration of factors such as request frequency and communication revenue. However, the above works are tailored for recommender system, which can not be directly migrated to the multi-modal domain field, mainly due to that they ignore the inherent complexity of multi-modal data. In this paper, we present CDC-MMPG, tailored for multi-modal on-device model adaptation via cloud-device collaboration, which simultaneously obtained effectiveness and efficiency.
@@
Domain Adaptation. Domain Adaptation (DA) plays a crucial role in transferring a network that was initially trained on a labeled source domain to effectively perform on a target domain. Various techniques have been developed to align the feature distributions between the source and target domains, such as maximum mean discrepancy and correlation alignment. These methods aim to reduce the distribution discrepancy and enable the network to generalize well across different domains, improving its performance in the target domain. Lots of methods are proposed to facilitate the development of the domain adaptation, such as multi-stage methods (Zhang et al., 2019; Chen et al., 2021), gradual transfer strategies (Chen et al., 2019), and so on. Recent advancements in domain adaptation embrace curriculum-based strategies inspired by curriculum learning to enhance the effectiveness of the adaptation process (Shu et al., 2019; Roy et al., 2021), Source-free domain adaptation (SFDA) has emerged as a response to privacy and copyright concerns, where access to the source domain data is restricted or unavailable (Zhang et al., 2019; Chen et al., 2021; Shu et al., 2019; Chen et al., 2019; Roy et al., 2021; Luo et al., 2020a; Gal et al., 2022; Ding et al., 2022; Wang et al., 2022; Yang et al., 2022; Sun et al., 2022; Litrico et al., 2023). SFDA presents a more challenging variant known as Test-Time Adaptation (TTA), which necessitates online adaptation during the inference phase. In addition, many researchers introduce the domain adaptation setting into different tasks to enhance model generalization (Wang et al., 2023b; Mosallanezhad et al., 2022; Ahmed et al., 2023; da Costa et al., 2022; Yang et al., 2023a; Ma et al., 2022; Wang et al., 2023c; Zhang et al., 2022; Lu et al., 2022; Zhao et al., 2022; Yang et al., 2023b; Saltori et al., 2023). In summary, domain adaptation (DA) methods focus on aligning feature distributions between the source and target domains. Recent advancements have highlighted the importance of curriculum-based strategies to improve the adaptation process. (Zhang et al., 2019; Chen et al., 2021; Shu et al., 2019; Chen et al., 2019; Roy et al., 2021; Ye et al., 2022). In this paper, we propose a model generation framework to solve the personalized adaption problem.

### v3

@@ 2. Related Works
Cloud-Device Collaboration. In the rapidly evolving sphere of cloud-device collaboration, deep learning has become a crucial component, strategically combining the features of cloud-based and on-device machine learning. Although Federated Learning (FL), notably represented by FedAVG (Collins et al., 2022), has traditionally been a significant player, its perceived simplicity has limited its applicability in various practical situations. This dynamic field has seen the rise of innovative methods and techniques designed to optimize the interaction between cloud resources and mobile devices. For instance, MPDA (Yan et al., 2022) skillfully utilizes cloud-based samples to enhance the performance of on-device models. Another interesting approach involves deploying multiple models with similar functionalities, coordinated by a Meta Controller for optimal task allocation, thus expanding the possibilities of collaborative model adaptation (Zhang et al., 2022b; Chen et al., 2023). Moreover, DUET (Lv et al., 2023b), inspired by the HyperNetwork concept (Yi et al., 2023; Xian et al., 2021; Zhang et al., 2020; Oscar Chang, 2020), simplifies the device model adaptation process, eliminating the need for on-device training (Zhong et al., 2024). IDEAL (Lv et al., 2023a) expands parameter generation-based models to the domain of recommender systems, particularly DC-CDR (Zhang et al., 2023b). This enhances the generalization capabilities of device recommendation models, though it requires consideration of factors such as request frequency and communication revenue. However, the above works are tailored for recommender system, which can not be directly migrated to the multi-modal domain field, mainly due to that they ignore the inherent complexity of multi-modal data. In this paper, we present CDC-MMPG, tailored for multi-modal on-device model adaptation via cloud-device collaboration, which simultaneously obtained effectiveness and efficiency.
@@
Domain Adaptation. Domain Adaptation (DA) plays a crucial role in transferring a network that was initially trained on a labeled source domain to effectively perform on a target domain. Various techniques have been developed to align the feature distributions between the source and target domains, such as maximum mean discrepancy and correlation alignment. These methods aim to reduce the distribution discrepancy and enable the network to generalize well across different domains, improving its performance in the target domain. Lots of methods are proposed to facilitate the development of the domain adaptation, such as multi-stage methods (Zhang et al., 2019; Chen et al., 2021; Li et al., 2024b; Ji et al., 2024; Li et al., 2023), gradual transfer strategies (Chen et al., 2019; Sun et al., 2022a; Li et al., 2024c; Tu et al., 2023), and so on. Recent advancements in domain adaptation embrace curriculum-based strategies inspired by curriculum learning to enhance the effectiveness of the adaptation process (Shu et al., 2019; Roy et al., 2021; Li et al., 2024a), Source-free domain adaptation (SFDA) has emerged as a response to privacy and copyright concerns, where access to the source domain data is restricted or unavailable (Zhang et al., 2019; Chen et al., 2021; Shu et al., 2019; Chen et al., 2019; Roy et al., 2021; Luo et al., 2020a; Gal et al., 2022; Ding et al., 2022; Wang et al., 2022; Yang et al., 2022; Sun et al., 2022b; Litrico et al., 2023). SFDA presents a more challenging variant known as Test-Time Adaptation (TTA), which necessitates online adaptation during the inference phase. In addition, many researchers introduce the domain adaptation setting into different tasks to enhance model generalization (Wang et al., 2023b; Mosallanezhad et al., 2022; Ahmed et al., 2023; da Costa et al., 2022; Yang et al., 2023a; Ma et al., 2022; Wang et al., 2023c; Zhang et al., 2022a; Lu et al., 2022; Zhao et al., 2022; Yang et al., 2023b; Saltori et al., 2023; Zhang et al., 2023a). In summary, domain adaptation (DA) methods focus on aligning feature distributions between the source and target domains. Recent advancements have highlighted the importance of curriculum-based strategies to improve the adaptation process. (Zhang et al., 2019; Chen et al., 2021; Shu et al., 2019; Chen et al., 2019; Roy et al., 2021; Ye et al., 2022). In this paper, we propose a model generation framework to solve the personalized adaption problem.